OpenStudy (anonymous):

10. Ann is moving from Houston to McKinney and rented a truck from U-Move truck rentals. The cost of a one-day truck rental is given by C(m) = 0.3m + 50 where m is the number of miles driven. If the cost of the truck rental is $131.00, how many miles did Ann drive the rental truck? 300 miles 230 miles 280 miles 270 miles

OpenStudy (anonymous):

131 = 0.3m + 50

OpenStudy (anonymous):

can you solve for m ?

OpenStudy (anonymous):

131=.3m+50 81=.3m 810=3m

OpenStudy (anonymous):

no

OpenStudy (anonymous):

subtract 50 from both sides 131-50 = 0.3m + 50 - 50 81 = 0.3m divide both sides by 0.3 m = 270